# Maximum Likelihood Estimation Based Complex-Valued Robust Chinese Remainder Theorem and Its Fast Algorithm

### v1

@@ Abstract
Recently, a multi-channel self-reset analog-to-digital converter (ADC) system with complex-valued moduli has been proposed. This system enables the recovery of high dynamic range complex-valued bandlimited signals at low sampling rates via the Chinese remainder theorem (CRT). In this paper, we investigate complex-valued CRT (C-CRT) with erroneous remainders, where the errors follow wrapped complex Gaussian distributions. Based on the existing real-valued CRT utilizing maximum likelihood estimation (MLE), we propose a fast MLE-based C-CRT (MLE C-CRT). The proposed algorithm requires only $2L$ searches to obtain the optimal estimate of the common remainder, where $L$ is the number of moduli. Once the common remainder is estimated, the complex number can be determined using the C-CRT. Furthermore, we obtain a necessary and sufficient condition for the fast MLE C-CRT to achieve robust estimation. Finally, we apply the proposed algorithm to ADCs. The results demonstrate that the proposed algorithm outperforms the existing methods.
@@
Chinese remainder theorem (CRT), real-valued CRT, complex-valued CRT (C-CRT), robust CRT, residue number system, analog-to-digital converter (ADC), self-reset ADC (SR-ADC).
@@ I Introduction
UNLIMITED sampling has garnered significant attention in recent years [1, 2, 3, 4, 5], where the traditional sampling is replaced by modulo sampling, which goes through the self-reset analog-to-digital converter (SR-ADC). Unlike the conventional sampling, unlimited sampling needs a sampling rate about 17 times faster than the Nyquist rate for the reconstruction of bandlimited signals. To avoid exceeding the Nyquist rate, multi-channel SR-ADC systems have been proposed in [6, 7, 8, 9], as shown in Fig. 1. In [7], the authors have proposed a complex-valued modulus multi-channel ADC architecture based on Gaussian integers, which offers a higher dynamic range compared to [6]. In [7], to recover complex-valued bandlimited signals, the moduli are classified into two types: Gaussian integers and positive integers. A signal value is then reconstructed using the two-stage robust Chinese remainder theorem (CRT) [10] for a real number. In the first stage, the remainders of the Gaussian integer moduli are used to recover the partial signal value through the closed-form robust CRT [11]. In the second stage, the remainders of the positive integer moduli are utilized to recover the complete signal value based on the first stage. It is demonstrated that the reconstruction is robust if the error conditions for both stages are satisfied. Furthermore, the SR-ADC with a complex-valued modulus has found applications in signal frequency estimation [12] and in massive multiple-input multiple-output (MIMO) systems [13].
@@
In addition to the above applications in ADC, CRT has many other applications in computer science, coding theory, and digital signal processing [14, 15]. However, the CRT is not robust as even a small error in any remainder may lead to a large error in the reconstruction. To overcome this shortcoming, a robust CRT has been studied in [16, 17, 18, 19, 20, 21] by utilizing remainder redundancy. The existing literature considers two types of remainder redundancy: 1) the remaining factors of the moduli after dividing by their greatest common divisor (gcd); and 2) the remaining factors of the moduli after dividing by their gcd are not pairwise coprime. For the first type of remainder redundancy, any two moduli have the same gcd. Furthermore, all the remainders modulo the gcd are identical and are referred to as the common remainder [22]. In [16], a searching-based method is proposed to address this redundancy. In [11], a closed-form CRT is introduced, assuming identical remainder error variances, thereby eliminating the need for search steps through a direct closed-form reconstruction process. In [21], a maximum likelihood estimation (MLE)-based algorithm is proposed, which optimally estimates the common remainder and the noises may have different variances. For the second type of remainder redundancy, there are at least two distinct groups of moduli, each having a different gcd [10, 20]. In [10], a multi-stage robust CRT method is proposed that enhances the robustness, with a potentially improved performance compared to the first type when the moduli are appropriately grouped. The robust CRT has numerous applications, such as in multi-channel SAR and InSAR systems [23, 24]. It has also been generalized for vectors [25, 26], for multiple integers [27, 28, 29, 30, 31], and for polynomials [32, 33].
The core of the robust CRT lies in limiting the reconstruction error level the same as that of the remainder errors. For real numbers, remainder errors (or noises) are typically measured using the circular distance based on the modulo operation [21]. However, the circular distance between two complex numbers involves both scaling and rotation in the complex plane. Hence, an error must be computed by considering both the real and imaginary parts simultaneously. Moreover, when the moduli are complex numbers, the reconstruction process becomes more complicated. When the complex moduli are pairwise conjugate each other, the reconstruction of a complex number can be solved by the two-stage robust CRT for real numbers [7]. To the best of our knowledge, there is no existing literature that addresses the reconstruction of a complex number from its erroneous remainders using a robust complex-valued CRT (C-CRT). In this paper, we focus on estimating an unknown complex number from its erroneous complex remainders, where the moduli are Gaussian integers, with their product and gcd being real-valued integers (or simply called integers). Additionally, the moduli are pairwise coprime after divided by their gcd [34]. Motivated by the work of [21], we propose a fast MLE-based algorithm for the robust C-CRT for complex numbers, which is more challenging compared to that for real numbers, particularly in the MLE model and the analysis of robust estimation. Our contributions are threefold. First, we propose an efficient algorithm to determine the MLE from erroneous remainders with wrapped complex Gaussian noises. Second, we provide the optimal estimate of the common remainder from complex erroneous remainders. The total number of the optimal common remainder candidates is $2L$ compared to $L$ in the real-valued case presented in [21], where $L$ is the number of moduli. Third, we derive a necessary and sufficient condition for the MLE-based C-CRT (MLE C-CRT) to be robust. Additionally, we apply the proposed fast MLE C-CRT algorithm to ADCs.
@@ II-A Basic Concepts for C-CRT
$\mathcal{F}_{\mathsf{M}}=\left\{\mathsf{M}(a+b\mathrm{i}):0\leq a,b<1\right\}.$ (2)
@@
For convenience, we denote $\left\langle\mathsf{N}\right\rangle_{\mathsf{M}}=\mathsf{r}$.
@@
Fig. 2 gives an illustration of the complex remainder set $\mathcal{F}_{\mathsf{M}}$ when $\mathsf{N}=2+5\mathrm{i}$. By (5), we have $\left\langle\mathsf{N}\right\rangle_{3+4\mathrm{i}}=-1+\mathrm{i}$ and $\left\langle\mathsf{N}\right\rangle_{4}=2+\mathrm{i}$. Clearly, if $\mathsf{M}$ is a real number, then $\langle\mathsf{N}\rangle_{\mathsf{M}}=\langle\mathrm{Re}(\mathsf{N})\rangle_{\mathsf{M}}+\mathrm{i}\langle\mathrm{Im}(\mathsf{N})\rangle_{\mathsf{M}}$. In this case, the modulo operation is performed separately on the real and imaginary parts of $\mathsf{N}$.
The C-CRT replaces moduli and remainders of the real-valued CRT with complex values. Now, we introduce the C-CRT when moduli are Gaussian integers. Unlike the CRT in rings, where the remainders must belong to an integral domain (see Theorem 17, Section 7.6 in [37]), this C-CRT allows the remainders to be any complex numbers, not just limited to Gaussian integers. For two Gaussian integers $\mathsf{\Gamma}_{i}$ and $\mathsf{\Gamma}_{j}$, we say that $\mathsf{\Gamma}_{i}$ and $\mathsf{\Gamma}_{j}$ are coprime if their common divisors are all ±$\pm 1$ and ±$\pm\mathrm{i}$. This coprimality is consistent with that for 2D integer matrices, i.e., if $\mathbf{\Gamma}_{i}$ and $\mathbf{\Gamma}_{j}$ are treated as two 2D integer matrices, then two Gaussian integers $\mathsf{\Gamma}_{i}$ and $\mathsf{\Gamma}_{j}$ are coprime if and only if their corresponding two 2D integer matrices $\mathbf{\Gamma}_{i}$ and $\mathbf{\Gamma}_{j}$ are coprime [25, 26, 36].
@@
Let $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{L}$ be pairwise coprime Gaussian integers with $|\mathsf{\Gamma}_{i}|\geq\sqrt{2}$ for all $i=1,2,\ldots,L$. Then, the system of congruences
$\mathsf{N}\equiv\mathsf{r}_{i}\mod\mathsf{\Gamma}_{i},\ i=1,2,\ldots,L,$ (6)
has a unique solution
$\mathsf{N}=\left\langle\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\mathsf{r}_{i}\right\rangle_{\mathsf{\Gamma}}$ (7)
in $\mathcal{F}_{\mathsf{\Gamma}}$, where $\mathsf{\Gamma}=\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$, $\mathsf{\gamma}_{i}=\frac{\mathsf{\Gamma}}{\mathsf{\Gamma}_{i}}$, and $\bar{\mathsf{\gamma}}_{i}$ is the modular multiplicative inverse of $\mathsf{\gamma}_{i}$ modulo $\mathsf{\Gamma}_{i}$, i.e., there exists a Gaussian integer $\bar{\mathsf{\Gamma}}_{i}$ such that
@@
First, we prove the existence. Let $\mathsf{q}=\left\lfloor\frac{\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\mathsf{r}_{i}}{\mathsf{\Gamma}}\right\rfloor$, by (7), we have $\mathsf{N}=\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\mathsf{r}_{i}-\mathsf{\Gamma}\mathsf{q}$. By (8), we have $\mathsf{N}=\sum_{i\neq k}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\mathsf{r}_{i}+\mathsf{(}1-\mathsf{\Gamma}_{k}\bar{\mathsf{\Gamma}}_{k})r_{k}-\mathsf{\Gamma}\mathsf{q}$ for $k=1,2,\ldots,L$. Hence, $\mathsf{N}\equiv\mathsf{r}_{k}\mod\mathsf{\Gamma}_{k}$.
@@
Remark 1: Different from the CRT in rings, Theorem 1 gives a reconstruction method for a complex number in $\mathcal{F}_{\mathsf{\Gamma}}$ not just Gaussian integer. In fact, the CRT in the ring of Gaussian integers can be easily generalized from that in the ring of integers. The C-CRT described here is for any complex number, which comes from the applications where the remainders are complex numbers.
Note that the general 2D-CRT for 2D vectors studied in [25] and [26] may not have the concise form in (7) similar to the conventional CRT for real integers. Another advantage of the above C-CRT over the general 2D-CRT is that it may be easier to find a set of pairwise coprime Gaussian integers [38] than that for 2D integer matrices. Furthermore, a necessary and sufficient condition was obtained in [39] for 2D integer matrices of the form (4) called skew-circulant matrices in [39] to be coprime.
We next consider the following system of congruences
$\mathsf{N}\equiv\mathsf{r}_{i}\mod M\mathsf{\Gamma}_{i},\ i=1,2,\ldots,L,$ (9)
where $\Gamma=\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$ and $M$ are integers. In other words, both the gcd and the least common multiple (lcm) of the moduli are integers and in this case $\mathcal{F}_{M\Gamma}$ is a square of sides on the real and imaginary axes.
By (9), for each $i$ we obtain that there exists a Gaussian integer $\mathsf{k}_{i}$ such that $\mathsf{r}_{i}-\mathsf{N}=M\mathsf{\Gamma}_{i}\mathsf{k}_{i}$. Thus, $\mathsf{r}_{i}\equiv\mathsf{N}\mod M$. Then, we have $\mathsf{r}_{i}\equiv\left\langle\mathsf{N}\right\rangle_{M}\mod M$. That is, all remainders modulo $M$ have the same value called the common remainder $\mathsf{r}^{c}$, which is in $\mathcal{F}_{M}$ and can be determined by
@@
Let
@@
Then, we have
@@
Hence, $\mathbf{\Gamma}_{1}^{-1}\mathbf{\Gamma}_{2}$ has two real eigenvalues, $\frac{b_{1}}{a_{1}}$ and $\frac{b_{2}}{a_{2}}$. This contradicts the fact that $\mathbf{\Gamma}_{1}^{-1}\mathbf{\Gamma}_{2}=\begin{pmatrix}-\frac{7}{25}&\frac{24}{25}\\ -\frac{24}{25}&-\frac{7}{25}\end{pmatrix}$ has two complex eigenvalues, $-\frac{7}{25}+\frac{24}{25}\mathrm{i}$ and $-\frac{7}{25}-\frac{24}{25}\mathrm{i}$.
@@
where $\left[z_{i}\right]$ satisfy $-\frac{1}{2}\leq z_{i}-\left[z_{i}\right]<\frac{1}{2}$ for $i=1,2$. It is easy to verify that if $\mathsf{r}=\mathsf{N}-\mathsf{M}\left[\frac{\mathsf{N}}{\mathsf{M}}\right]$, then $|\mathsf{r}|<|\mathsf{M}|$. Hence, we can recursively obtain the modular inverse of the Gaussian integer using the Euclidean algorithm. For example, we let $\mathsf{n}=19+8\mathrm{i}$ and $\mathsf{m}=3+4\mathrm{i}$. Clearly, $\left[\frac{\mathsf{n}}{\mathsf{m}}\right]=4-2\mathrm{i}$. Hence, $\mathsf{n}=(4-2\mathrm{i})\mathsf{m}+(-1-2\mathrm{i})$. Since $\left[\frac{\mathsf{m}}{-1-2\mathrm{i}}\right]=-2$, we have $\mathsf{m}=-2(-1-2\mathrm{i})+1$. Then, we have the following Bezout’s identity:
@@
That is, $(-7+4\mathrm{i})\mathsf{m}\equiv 1\mod\mathsf{n}$. In other words, the modular multiplicative inverse of $\mathsf{m}$ modulo $\mathsf{n}$ is $-7+4\mathrm{i}$.

### v3

@@ Abstract
In this paper, we investigate complex-valued Chinese remainder theorem (C-CRT) with erroneous remainders, where the moduli are Gaussian integers and the errors follow wrapped complex Gaussian distributions. Based on the existing real-valued CRT utilizing maximum likelihood estimation (MLE), we propose a fast MLE-based C-CRT (MLE C-CRT). The proposed algorithm requires only $2L$ searches to obtain the optimal estimate of the common remainder, where $L$ is the number of moduli. Once the common remainder is estimated, the complex number can be determined using the C-CRT. Furthermore, we obtain a necessary and sufficient condition for the fast MLE C-CRT to achieve robust estimation. Finally, we apply the proposed algorithm to a multi-channel self-reset analog-to-digital converter (ADC) system with Gaussian integers as moduli, which enables the recovery of high dynamic range complex-valued bandlimited signals at the Nyquist sampling rate. The results demonstrate that the proposed algorithm outperforms the existing methods.
@@
Chinese remainder theorem (CRT), real-valued CRT, complex-valued CRT (C-CRT), robust CRT, residue number system, multi-channel self-reset (SR) analog-to-digital converter (ADC).
@@ I Introduction
The Chinese remainder theorem (CRT) is a fundamental theorem in ring theory, widely applied in computer science, coding theory, and digital signal processing [1, 2]. However, the CRT is not robust as even a small error in any remainder may lead to a large error in the reconstruction. To overcome this shortcoming, a robust CRT has been studied in [3, 4, 5, 6, 7, 8, 9, 10] by utilizing remainder redundancy. The existing literature mainly considers two types of remainder redundancy: 1) the remaining factors of the moduli after being divided by their greatest common divisor (gcd) are pairwise coprime; and 2) the remaining factors of the moduli after being divided by their gcd are not pairwise coprime. For the first type of remainder redundancy, any two moduli have the same gcd. Furthermore, all the remainders modulo the gcd are identical and are referred to as the common remainder [11]. In [3], a searching-based method is proposed to address this redundancy. In [7], a closed-form CRT is introduced, assuming identical remainder error variances, thereby eliminating the need for search steps through a direct closed-form reconstruction process. In [10], a maximum likelihood estimation (MLE)-based algorithm is proposed, which optimally estimates the common remainder and the noises may have different variances. For the second type of remainder redundancy, there are at least two distinct groups of moduli, each having a different gcd [8, 9]. In [9], a multi-stage robust CRT method is proposed that enhances the robustness, with a potentially improved performance compared to the first type when the moduli are appropriately grouped. The robust CRT has numerous applications, such as in multi-channel SAR and InSAR systems [12, 13]. It has also been generalized for vectors [14, 15], for multiple integers [16, 17, 18, 19, 20], and for polynomials [21, 22].
In this paper, we propose robust complex-valued CRT (C-CRT) with Gaussian integers as moduli to robustly determine a complex number from its remainders modulo several Gaussian integers. It can be thought of as a generalization of the robust CRT for real numbers in [7, 10]. Note that the robust CRT for real numbers is able to have the reconstruction error level the same as that of the remainder errors (or noises) that are typically measured using the circular distance based on the modulo operation [10]. However, the circular distance between two complex numbers involves both scaling and rotation in the complex plane. Hence, an error must be computed by considering both the real and imaginary parts simultaneously. Moreover, when the moduli are complex numbers, the reconstruction process becomes more complicated. When the complex moduli are pairwise conjugate each other, the reconstruction of a complex number can be solved by the two-stage robust CRT for real numbers [23]. To the best of our knowledge, there is no robust C-CRT that addresses the reconstruction of a complex number from its erroneous remainders directly. In this paper, we propose a robust C-CRT with Gaussian integers as moduli, where the product and the gcd of the moduli are real-valued integers (or simply called integers). Additionally, the moduli are pairwise coprime after divided by their gcd. Motivated by the work of [10], we propose a fast MLE-based algorithm for the robust C-CRT for complex numbers, which is more challenging compared to that for real numbers, particularly in the MLE model and the analysis of robust estimation.
@@
The proposed fast MLE C-CRT algorithm can be applied in the multi-channel self-reset analog-to-digital converter (SR-ADC) system proposed in [25, 23, 26, 27], as shown in Fig. 1. Compared to the single-channel ADC system, it offers a higher dynamic range and is capable of reconstructing bandlimited signals with sampling at the Nyquist rate. Unlike the traditional ADC-based continuous-time signal recovery, this system first reconstructs the sampled values from their multiple modulo samples before recovering a continuous-time signal. In [23], the authors have proposed a complex-valued modulus multi-channel ADC architecture based on Gaussian integers, which offers a higher dynamic range compared to [25]. To recover sampled values of a complex-valued bandlimited signal, the moduli are classified into two types: Gaussian integers and positive integers. A sampled value is then reconstructed using the two-stage robust CRT [9] for a real number. In the first stage, the remainders of the Gaussian integer moduli are used to recover the partial signal value through the closed-form robust CRT [7]. In the second stage, the remainders of the positive integer moduli are utilized to recover the complete signal value based on the first stage. It is demonstrated that the reconstruction is robust if the error conditions for both stages are satisfied.
Our contributions are fourfold. First, we propose an efficient algorithm for the C-CRT to determine the MLE from erroneous remainders with wrapped complex Gaussian noises. Second, we provide the optimal estimate of the common remainder from complex erroneous remainders. The total number of the optimal common remainder candidates is $2L$ compared to $L$ in the real-valued case presented in [10], where $L$ is the number of moduli. Third, we derive a necessary and sufficient condition for the MLE-based C-CRT (MLE C-CRT) to be robust. Forth, we apply our proposed robust C-CRT to multi-channel SR-ADCs with improved performance.
@@ II-A Basic Concepts for C-CRT
$\mathcal{F}_{\mathsf{M}}=\left\{\mathsf{M}(a+b\mathrm{i}):0\leq a,b<1\right\},$ (2)
and $\mathsf{r}$ is called the remainder of $\mathsf{N}$ modulo $\mathsf{M}$.
@@
For convenience, we denote $\left\langle\mathsf{N}\right\rangle_{\mathsf{M}}=\mathsf{r}=\mathsf{N}\mod\mathsf{M}$.
@@
Fig. 2 gives an illustration of the complex remainder set $\mathcal{F}_{\mathsf{M}}$ when $\mathsf{N}=2+5\mathrm{i}$. By (5), we have $\left\langle\mathsf{N}\right\rangle_{3+4\mathrm{i}}=-1+\mathrm{i}$ and $\left\langle\mathsf{N}\right\rangle_{4}=2+\mathrm{i}$. Clearly, if $\mathsf{M}$ is a real number, then
$\langle\mathsf{N}\rangle_{\mathsf{M}}=\langle\mathrm{Re}(\mathsf{N})\rangle_{\mathsf{M}}+\mathrm{i}\langle\mathrm{Im}(\mathsf{N})\rangle_{\mathsf{M}}.$
In this case, the modulo operation is performed separately on the real and imaginary parts of $\mathsf{N}$.
The C-CRT replaces moduli and remainders of the real-valued CRT with complex values. Now, we introduce the C-CRT when moduli are Gaussian integers. Unlike the CRT in rings, where the remainders must belong to an integral domain (see Theorem 17, Section 7.6 in [29]), this C-CRT allows the remainders to be any complex numbers, not just limited to Gaussian integers similar to the CRT for real numbers studied in [7, 10]. The problem is as follows. Let $\mathsf{\Gamma}_{i}$, $i=1,2,\ldots,L$, be $L$ Gaussian integers as moduli, $\mathsf{N}$ be a complex number, and $\mathsf{r}_{i}=\mathsf{N}\mod\mathsf{\Gamma}_{i}$, $i=1,2,\ldots,L$, be its remainders, where $\mathsf{N}$ and $\mathsf{r}_{i}$, $i=1,2,\ldots,L$, may not necessarily be Gaussian integers as explained above. Thus, we have the following system of congruences:
$\mathsf{N}=\mathsf{k}_{i}\mathsf{\Gamma}_{i}+\mathsf{r}_{i},\ i=1,2,\ldots,L,$ (6)
where $\mathsf{k}_{i}$, $i=1,2,\ldots,L$, are unknown Gaussian integers called folding Gaussian integers. The problem is to determine $\mathsf{N}$ from its remainders $\mathsf{r}_{i}$, $i=1,2,\ldots,L$. This problem occurs in multi-channel SR-ADC for complex-valued bandlimited signals [23].
We next present C-CRT. For two Gaussian integers $\mathsf{\Gamma}_{i}$ and $\mathsf{\Gamma}_{j}$, we say that $\mathsf{\Gamma}_{i}$ and $\mathsf{\Gamma}_{j}$ are coprime if their common divisors are all $\pm 1$ and $\pm\mathrm{i}$. This coprimality is consistent with that for 2D integer matrices, i.e., if $\mathbf{\Gamma}_{i}$ and $\mathbf{\Gamma}_{j}$ are treated as two 2D integer matrices, then two Gaussian integers $\mathsf{\Gamma}_{i}$ and $\mathsf{\Gamma}_{j}$ are coprime if and only if their corresponding two 2D integer matrices $\mathbf{\Gamma}_{i}$ and $\mathbf{\Gamma}_{j}$ are coprime [30].
@@
Let $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{L}$ be pairwise coprime Gaussian integers with $|\mathsf{\Gamma}_{i}|\geq\sqrt{2}$ for $i=1,2,\ldots,L$. Then, for a complex number $\mathsf{N}\in\mathcal{F}_{\mathsf{\Gamma}}$, the system of congruences (6) has a unique solution
$\mathsf{N}=\left\langle\mathsf{r}_{1}-\left\lfloor\mathsf{r}_{1}\right\rfloor+\sum_{i=1}^{L}\bar{\gamma}_{i}\mathsf{\gamma}_{i}\left\lfloor\mathsf{r}_{i}\right\rfloor\right\rangle_{\mathsf{\Gamma}},$ (7)
where $\mathsf{\Gamma}=\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$, $\mathsf{\gamma}_{i}=\frac{\mathsf{\Gamma}}{\mathsf{\Gamma}_{i}}$, and $\bar{\mathsf{\gamma}}_{i}$ is the modular multiplicative inverse of $\mathsf{\gamma}_{i}$ modulo $\mathsf{\Gamma}_{i}$, i.e., there exists a Gaussian integer $\bar{\mathsf{\Gamma}}_{i}$ such that
@@
Let $\mathsf{N}^{\prime}=\mathsf{N}-\left\lfloor\mathsf{N}\right\rfloor$ and $\mathsf{r}_{i}^{\prime}=\mathsf{r}_{i}-\left\lfloor\mathsf{r}_{i}\right\rfloor$ for $i=1,2,\ldots,L$. By (6), we have
$\mathsf{N}^{\prime}-\mathsf{r}_{i}^{\prime}=\mathsf{k}_{i}\mathsf{\Gamma}_{i}+\left\lfloor\mathsf{r}_{i}\right\rfloor-\left\lfloor\mathsf{N}\right\rfloor\in\mathbb{Z}[\mathrm{i}],\ i=1,2,\ldots,L.$ (9)
Since $\mathsf{N}^{\prime}\in\mathcal{F}_{1}$ and $\mathsf{r}_{i}^{\prime}\in\mathcal{F}_{1}$. We have $\mathrm{Re}(\mathsf{N}^{\prime}-\mathsf{r}_{i}^{\prime})\in(-1,1)$ and $\mathrm{Im}(\mathsf{N}^{\prime}-\mathsf{r}_{i}^{\prime})\in(-1,1)$. Then, we obtain from (9) that $\mathsf{N}^{\prime}=\mathsf{r}_{i}^{\prime}$ holds for each $i$. Hence,
$\mathsf{r}_{1}^{\prime}=\mathsf{r}_{2}^{\prime}=\cdots=\mathsf{r}_{L}^{\prime}.$
Consequently, (6) can be rewritten as
$\mathsf{N}-\mathsf{r}_{1}^{\prime}=\mathsf{k}_{i}\mathsf{\Gamma}_{i}+\left\lfloor\mathsf{r}_{i}\right\rfloor,\ i=1,2,\ldots,L.$
Since $\mathsf{\Gamma}_{1},\mathsf{\Gamma}_{2},\ldots,\mathsf{\Gamma}_{L}$ are pairwise coprime, according to the CRT over rings,
$\mathsf{N}-\mathsf{r}_{1}^{\prime}\equiv\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\left\lfloor\mathsf{r}_{i}\right\rfloor\mod\mathsf{\Gamma}.$
That is, there exists a Gaussian integer $\mathsf{k}$ such that
$\mathsf{N}-\mathsf{r}_{1}^{\prime}-\sum_{i=1}^{L}\bar{\mathsf{\gamma}}_{i}\mathsf{\gamma}_{i}\left\lfloor\mathsf{r}_{i}\right\rfloor=\mathsf{k}\mathsf{\Gamma}.$
Therefore, (7) is a solution of (6) in $\mathcal{F}_{\mathsf{\Gamma}}$.
@@
Remark 1: Different from the CRT in rings, Theorem 1 gives a reconstruction method for a complex number in $\mathcal{F}_{\mathsf{\Gamma}}$ (not just a Gaussian integer) from its remainders. In fact, the CRT in the ring of Gaussian integers can be easily generalized from that in the ring of integers. The C-CRT described here is for any complex number, which comes from the applications where the unknown $\mathsf{N}$ and its remainders are complex numbers.
Note that the general 2D-CRT for 2D vectors studied in [14] and [15] may not have the concise form in (7) similar to the conventional CRT for real integers. Another advantage of the above C-CRT over the general 2D-CRT is that it may be more convenient to find a set of pairwise coprime Gaussian integers [31] than that for 2D integer matrices. Furthermore, a necessary and sufficient condition was obtained in [32] for 2D integer matrices of the form (4) called skew-circulant matrices in [32] to be coprime.
As explained in Introduction, the key for the robust CRT for real numbers in the literature is to have some redundancies in the remainders. One of such redundancies is to have a non-unit gcd among all the moduli. We next consider the moduli of the forms $M\mathsf{\Gamma}_{i}$ where $M$ is the gcd of all the moduli. Thus, we have the following system of congruences
$\mathsf{N}=\mathsf{k}_{i}M\mathsf{\Gamma}_{i}+\mathsf{r}_{i},\ i=1,2,\ldots,L,$ (10)
where $\Gamma=\prod_{i=1}^{L}\mathsf{\Gamma}_{i}$ and $M$ are both assumed positive integers, $\mathsf{k}_{i}$, $i=1,2,\ldots,L$, are the unknown folding Gaussian integers. In other words, both the gcd and the least common multiple (lcm) of the moduli are assumed integers and in this case $\mathcal{F}_{M\Gamma}$ is a square of sides on the real and imaginary axes. In the following we generalize Theorem 1 to solve (10). By (10), we have
$\mathsf{r}_{i}\equiv\mathsf{N}\mod M.$
Then,
$\mathsf{r}_{i}\equiv\left\langle\mathsf{N}\right\rangle_{M}\mod M.$
That is, all remainders modulo $M$ have the same value called the common remainder $\mathsf{r}^{c}$, which is in $\mathcal{F}_{M}$ and can be determined by
@@
It follows that $\mathsf{r}_{i}-\mathsf{r}^{c}\in M\mathbb{Z}[\mathrm{i}]$. Let
@@
Then, we have $\mathsf{q}_{i}\in\mathbb{Z}[\mathrm{i}]$ and
@@
Hence, $\mathbf{\Gamma}_{1}^{-1}\mathbf{\Gamma}_{2}$ has two real eigenvalues, $\frac{b_{1}}{a_{1}}$ and $\frac{b_{2}}{a_{2}}$. This contradicts the fact that $\mathbf{\Gamma}_{1}^{-1}\mathbf{\Gamma}_{2}=\begin{pmatrix}-\frac{7}{25}&\frac{24}{25}\\ -\frac{24}{25}&-\frac{7}{25}\end{pmatrix}$ has two complex eigenvalues, $\frac{7}{25}+\frac{24}{25}\mathrm{i}$ and $\frac{7}{25}-\frac{24}{25}\mathrm{i}$.
@@
where $\left[z_{i}\right]$ satisfy $\frac{1}{2}\leq z_{i}-\left[z_{i}\right]<\frac{1}{2}$ for $i=1,2$. It is easy to verify that if $\mathsf{r}=\mathsf{N}-\mathsf{M}\left[\frac{\mathsf{N}}{\mathsf{M}}\right]$, then $|\mathsf{r}|<|\mathsf{M}|$. Hence, we can recursively obtain the modular inverse of the Gaussian integer using the Euclidean algorithm. For example, we let $\mathsf{n}=19+8\mathrm{i}$ and $\mathsf{m}=3+4\mathrm{i}$. Clearly, $\left[\frac{\mathsf{n}}{\mathsf{m}}\right]=4-2\mathrm{i}$. Hence, $\mathsf{n}=(4-2\mathrm{i})\mathsf{m}+(-1-2\mathrm{i})$. Since $\left[\frac{\mathsf{m}}{-1-2\mathrm{i}}\right]=-2$, we have $\mathsf{m}=-2(-1-2\mathrm{i})+1$. Then, we have the following Bezout’s identity:
@@
Hence, the modular multiplicative inverse of $\mathsf{m}$ modulo $\mathsf{n}$ is $7+4\mathrm{i}$.